OpenStudy (anonymous):

A ladder 26ft long leans against a vertical wall. The lower end is being moved away from the wall at a rate of 5ft/sec how fast is the top decreasing when the lower end is 10 ft. from the wall

OpenStudy (anonymous):

x² + y² = z² -> y = √ ( 26² - 10² ) = 24 2xdx/dt + 2ydy/dt = 0 => dy/dt = - x dx/dt * ( 1/y) = - 5 * 10 / 24 = - 2.08 ft/sec